OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A.) 5x - 8y = 9 12x - 5y = 7 B.) 3x + 6y = 8 y = 3x - 8 C.) y = 2x - 7 x + 3y = 3 D.) 2y = -3x + 5 2x - 3y = 4

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

I'm putting them in my grapher thingy right now

OpenStudy (anonymous):

Thanksss

OpenStudy (anonymous):

You're welcome. The answer 'tis D

OpenStudy (anonymous):

thanks! can you answer more questions please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Given the linear equation 3x + y = 5, find the slope of its graph. -3 -5 3 5

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Given the linear equation 3y + 4x =12, find the y-intercept. Option A: -4/3 Option B: -4 Option C: 4/3 Option D: 4

OpenStudy (anonymous):

y-intercept is 4

OpenStudy (anonymous):

Write an equation in slope-intercept form for the line passing through (-1, 5) and (5, 8). Option A: y= -1/2x +11/2 Option B: y= -1/2x - 11/2 Option C: y=1/2x - 11/2 Option D: y=1/2x +11/2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

YAY!!! Thanks for the fan!!! =) (= <3

OpenStudy (anonymous):

Solve the system of equations. 3x + 3y = -10   x + 3y = 0 (Points : 5)        Option A: x = -3/2; y = -1/2        Option B: x = 3; y = 1       Option C: x =-5; y = 5/3        Option D: x = 3; y = -5 no problem!! and thanks for helping~

OpenStudy (anonymous):

'Tis C. And you're welcome

OpenStudy (anonymous):

Solve for x. x - 1/12 = 7/2 (Points : 5) 3 Option A: x= 4- (the fraction is supposed to be 4 and 3 7ths) 7 Option B: x = 6 Option C: x = 41 Option D: x = 43

OpenStudy (anonymous):

this is the last question

OpenStudy (anonymous):

wait i typed the question wrong

OpenStudy (anonymous):

let me re-type it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh wait never mind its right ahaha

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that's weird none of the answers seem to be right

OpenStudy (anonymous):

let me check again

OpenStudy (anonymous):

never mind its 43

OpenStudy (anonymous):

i typed it in wrong at first

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

wait which one was it?

OpenStudy (anonymous):

It's x=43 in other words its Option D

OpenStudy (anonymous):

ohh ok, thank you!!

OpenStudy (anonymous):

i got 12/14 right so thanks!!

OpenStudy (anonymous):

you're welcome!!! =) (= <3

OpenStudy (anonymous):

Happy to help!!!

OpenStudy (anonymous):

if you want i have other questions, if you don't its ok!

OpenStudy (anonymous):

I can answer more that's fine

OpenStudy (anonymous):

thankyou!!

OpenStudy (anonymous):

It's A. (-1,7)

OpenStudy (anonymous):

Also known as the intersection point

OpenStudy (anonymous):

second one!

OpenStudy (anonymous):

it's the intersect (where they meet)

OpenStudy (anonymous):

so is it C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about this?

OpenStudy (anonymous):

Infinitely many. That's because they are the same line.

OpenStudy (anonymous):

this?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

these are the last questions of the quiz

OpenStudy (anonymous):

5. (4,0) 6. Is no solution (the lines are parallel)

OpenStudy (anonymous):

i got 6/6 thanks!!

OpenStudy (anonymous):

Sorry for the wait I had to take my dogs out (I have 5 of them and two fight so we have to keep them separated)

OpenStudy (anonymous):

YAY!!!

OpenStudy (anonymous):

wow 5 dogs haha

OpenStudy (anonymous):

I like 100%'s

OpenStudy (anonymous):

Note to EVERYONE: I'm VERY random sometimes

OpenStudy (anonymous):

haha me too, i wasn't doing so good so i had to take these online tests to bring up my grade and i have alot to do -__- lol

OpenStudy (anonymous):

Yeah but I love them all equally and I wouldn't traded them for anything (Not even the Akatsuki from Naruto Shippuden)

OpenStudy (anonymous):

i used to have a chinese pug

OpenStudy (anonymous):

that's why I'm doing Florida Virtual School. Well actually I had and have very good grades but I was bored in school so yeah

OpenStudy (anonymous):

OH!!! I like pugs they're so cute

OpenStudy (anonymous):

I have 3 Boxers and 2 English bulldogs (they're ALL girls) (Dad doesn't like boy animals)

OpenStudy (anonymous):

my grades were getting low because i went to florida for vacation and i had a lot of catching up to do! haha and yeah i do an ohio virtual school

OpenStudy (anonymous):

my mom doesn't like dogs at all lol

OpenStudy (anonymous):

LOL!!! I Live in Florida!!! That's kinda funny. and ironic

OpenStudy (anonymous):

haha yeah!

OpenStudy (anonymous):

But seriously it's ACTUALLY called Florida Virtual School (FLVS)

OpenStudy (anonymous):

mines called Ohio Virtual Academy (OHVA) from k12

OpenStudy (anonymous):

These people come up with such non-creative names don't they

OpenStudy (anonymous):

hahaha yes they do

OpenStudy (anonymous):

You know something. I've been doing Physics while helping you. I got bored

OpenStudy (anonymous):

That's what I do when I get bored with school. I go on here and see if I can help anyone

OpenStudy (anonymous):

well thats nice of you! like i have sooooo much to do because it's the end of the semester! i just want the semester to end already!

OpenStudy (anonymous):

Which Semester are you in 1 or 2? I'm in semester 2 of Physics. Actually getting ready to do the Chapter test for the last chapter an then all I have left is the Semester exam

OpenStudy (anonymous):

right now its the end of the 1st semester. the 2nd semester starts on the 18th

OpenStudy (anonymous):

Oh well um. You can move at your own pace in FLVS and so I've already finished semester 1 (I did that before Christmas)

OpenStudy (anonymous):

And thankfully there are only 3 chapter in semester 1 and 4 chapters in semester 2 so it's not that hard really

OpenStudy (anonymous):

wow that sounds much easier because here we don't move on our own pace and we have like 8 or 9 chapters per semester

OpenStudy (anonymous):

Yeah, you see the entire reason that I started homeschooling was because I had to do extra homework because SOMEONE ELSE didn't understand. Gosh I hated have to do more work cause of someone else. Not that it's their fault that they don't understand but I just think that it's unfair to give the kids that know the stuff extra work as well.

OpenStudy (anonymous):

yeah that isn't fair and i started homeschooling because the schools here are really bad

OpenStudy (anonymous):

Oh that sucks

OpenStudy (anonymous):

Hey I got to go take my chapter 7 test for Physics. I'm not logging out in case I need to ask a question but I'm gonna be technically off for a bit okay.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Which equation is shown on the graph? A. y = x – 3 B. y = 2x – 3 C. y = x + 3 D. y = 3x